# Identifying Aspects in Peer Reviews

### v2

@@ Abstract
Peer review is central to academic publishing, but the growing volume of submissions is straining the process. This motivates the development of computational approaches to support peer review. While each review is tailored to a specific paper, reviewers often make assessments according to certain aspects such as Novelty, which reflect the values of the research community. This alignment creates opportunities for standardizing the reviewing process, improving quality control, and enabling computational support. While prior work has demonstrated the potential of aspect analysis for peer review assistance, the notion of aspect remains poorly formalized. Existing approaches often derive aspects from review forms and guidelines, yet data-driven methods for aspect identification are underexplored. To address this gap, our work takes a bottom-up approach: we propose an operational definition of aspect and develop a data-driven schema for deriving aspects from a corpus of peer reviews. We introduce a dataset of peer reviews augmented with aspects and show how it can be used for community-level review analysis. We further show how the choice of aspects can impact downstream applications, such as LLM-generated review detection. Our results lay a foundation for a principled and data-driven investigation of review aspects, and pave the path for new applications of NLP to support peer review. [^1] Our code and data are available at https://anonymous.4open.science/r/aspects-in-reviews-737E/.
@@ 1 Introduction
Peer review is an essential part of academic publishing. It is a complex, multifaceted task that requires a range of competencies including paper understanding, domain-specific knowledge, and critical thinking Shah (2022); Yuan et al. (2022). Ensuring review quality, especially among novice reviewers, is an open challenge Stelmakh et al. (2021a, b); Sun et al. (2024a). The increasing volume of publications puts further strain on the process, which motivates the development of computational approaches to support different stages of peer review, from reading the paper to the final decision-making by the program committees Arous et al. (2021); Checco et al. (2021); Stelmakh et al. (2021a); Shah (2022); Schulz et al. (2022); Yuan et al. (2022); Lin et al. (2023b); Kuznetsov et al. (2024).
Peer reviews are the central component of the reviewing process. While individual reviews can vary widely, reviewers within the same community tend to focus on a specific set of general quality categories, or aspects, such as Clarity and Novelty. These aspects can be found in review forms, instructional materials, guidelines, and the resulting review texts. Aspects allow comparison of submissions across different dimensions of quality. A comprehensive set of aspects shared among reviewers is critical for ensuring review quality and consistency, and prior work has demonstrated the potential of aspect-based tools to support the review writing process Sun et al. (2024a, b).
Yet, several open questions remain. First, what is an aspect? Most prior work derives aspects from review forms and guidelines. The lack of an operational definition of aspect prevents the comparison of aspect schemata across studies. Second, what aspect granularity is appropriate for different tasks? While prior work operates with top-down, coarse-grained aspect schemata derived from review forms and guidelines, it is unclear whether they are comprehensive or provide sufficient granularity for NLP applications. Third, how can fine-grained aspect analysis support peer review? While aspects have been applied to certain tasks, the tasks that require a higher level of granularity are underexplored.
@@
Our work paves the path for data-driven analysis of aspects and enables new NLP applications to support peer review. Our method offers a bottom-up perspective that complements existing top-down schemata, and can be integrated with them by, for example, allowing domain experts to refine LLM-derived aspects into higher-level, domain-specific categories. Together, these perspectives enable a more comprehensive aspect-based analysis in peer review. To summarize, this work contributes the following:
@@ 2.2 Aspect in peer review
Beyond their role in review forms and guidelines, aspects are utilized in several other contexts related to peer review. Aspects are used in analyzing the sentiment of reviews and discovering the discourse relations within reviews Chakraborty et al. (2020); Kennard et al. (2022). In automatic review generation, Wang et al. (2020) generate reviews using aspect-based knowledge graphs, and Gao et al. (2024) prompt LLMs with aspect-based questions to generate reviews. Aspects are included in supporting systems for review writing, which have been shown to improve the comprehensiveness of the reviews written by reviewers with varying levels of experience Sun et al. (2024a, b). Table 1 summarizes the aspects used in these studies.
@@ 2.3 Quality of review writing
Defining the desiderata for review writing is essential to assess its effectiveness Jefferson et al. (2002). However, the desiderata are often not well defined or operationalizable in terms of automatic measurement Kuznetsov et al. (2024). While using simple proxies such as “helpfulness” may seem like a straightforward way to assess review quality, research has revealed that such kinds of evaluation can be biased in several ways – for example, an evaluator may be biased towards longer reviews, or those that recommend acceptance of their papers Wang et al. (2021); Goldberg et al. (2023).
@@ 2.4 LLM-generated review detection
The strong capabilities of state-of-the-art LLMs in text generation have led to the need for detectors to identify LLM-generated contents and to prevent potential misuse Clark et al. (2021); Gao et al. (2023); Wu et al. (2023). In peer review, LLMs pose a risk that reviewers may exploit LLMs to produce reviews entirely, which raises serious ethical concerns Yu et al. (2024). Current strategies for detecting LLM-generated reviews align with methods for detecting LLM-generated texts: (a) LLM-as-a-judge, which prompts LLMs to identify LLM-generated contents Zheng et al. (2023), (b) detection models, which are fine-tuned on both human and LLM-generated texts Guo et al. (2023), (c) reference-based methods, which compare the similarity between a candidate text and one generated by an LLM Gehrmann et al. (2019); Ippolito et al. (2020); Liang et al. (2024a); Yu et al. (2024). In our work, we provide an alternative approach – we demonstrate a comprehensive, fine-grained aspect set helps the detection of LLM-generated reviews.
@@ 3.2 Method
To establish a comprehensive set of aspects, we selected paper reviews from NLP and machine learning (ML) conferences across different time periods. We randomly selected 50 papers from each of the NLPeer Dycke et al. (2023) and EMNLP23222Publicly available through the OpenReview API. datasets. We used the keywords “natural” and “language” to filter NLP-related papers from ICLR. We randomly selected 50 NLP-related papers from each of the ICLR conferences from 2020 to 2024. We selected 350 papers in total, corresponding to 1094 reviews. We segmented the reviews into sentences using NLTK Punkt Tokenizer Bird and Loper (2004) and performed the identification at the sentence level. See Appendix B for more details.
@@
We notice that our post-processing method may lead to unrelated terms being grouped together. For example, using Improvement as a keyword groups Performance Improvement and Improvement Recommendation (i.e., suggestions to improve paper quality) together. We manually checked the results to verify and correct inappropriate groupings.
@@ 3.4 Validity check
Since LLMs have been shown to be sensitive to prompts Chen et al. (2023); Ajith et al. (2024), we experimented with different prompts and temperature to assess the consistency of the model annotations. We used three consistency metrics: exact match, BERTScore similarity Zhang et al. (2020), and Jaccard similarity. We show in Table 10 that the consistency between annotations generated under different settings is moderate to strong. 45.50% and 67.14% of the annotations obtained using different prompts and temperature have BERTScore similarities greater than 0.9. The increase in exact matches between the raw annotations and those mapped to the coarse label set suggests that differences introduced by varying prompts and temperature do not significantly affect the categorization, as most of them still fall within the same label category. The Jaccard similarities of aspects mapped to the coarse label set also indicate strong consistency. Table 11 and 12 help interpret these scores by showing examples of how pairs of texts and sets correspond to different BERTScore and Jaccard similarities.
@@ 4 Aspect prediction
$f:\begin{cases}\text{PAP, }p\rightarrow\hat{A_{p}};\\ \text{RAP, }r\rightarrow\hat{A_{r}};\end{cases}$ (1)
@@ 4.1 Method
RAP differs from the task described in Section 3 in that it is implemented using a supervised approach. We utilized our curated data to train models to identify aspects within reviews. We segmented the reviews into sentences using NLTK.
@@ 4.2 Results
In general, models perform better on RAP and the coarse label set yields higher performance across both tasks.
@@ 5.2 Review comparison
We obverve that LLM-generated reviews show a higher degree of similarity to each other in terms of aspects. This suggests that LLM-generated reviews may be more generic than human-written ones, with the model tending to comment on similar sets of aspects across different papers. This could also imply that the quality of LLM-generated reviews is still lacking in terms of specificity.
@@ 5.3 LLM-generated review detection
$t=\begin{cases}\text{intra, if }r,r_{i}\in R_{p},\,p\in P,\,r\neq r_{i};\\ \text{inter, if }r\in R_{p},\,r_{i}\in R_{q},\,p,q\in P,\,p\neq q;\end{cases}$
@@ 5.4 Recommendations
Based on our results, we make the following recommendations on selecting aspect granularity. The coarse label set is more suitable for high-level analysis tasks, such as analyzing review focus across different tracks and venues (Section 5.1). The fine label set is more appropriate for tasks that require nuanced analysis, where capturing specific and detailed feedback is critical, such as review comparison (Section 5.2) and LLM-generated review detection (Section 5.3). In some cases, a hybrid strategy might be the best option. For example, when evaluating review quality, one can first apply the coarse label set to assess coverage (i.e., whether key aspects are addressed), and then use the fine label set to assess specificity of the review. Aspect granularity is a design choice, and the optimal configuration depends on the task and user needs.
@@ 6 Conclusion
For human annotation, we recognize that using aspect labels from our taxonomy may introduce potential label bias. A straightforward solution to this would be to ask the annotators to do an open-ended annotation task without pre-defined labels. However, such study is very challenging in terms of the cognitive burden on annotators and ensuring annotation consistency. Automation bias is also a concern. While we address this by asking annotators to provide explanations for their annotations, follow-up work can explore alternative experimental strategies to measure and mitigate automation bias, e.g., through the use of distractor items.
@@
Finally, in Section 5.3, we only focus on the commonly used end-to-end general prompts for our review generator. We did not consider prompts that involve paper-specific aspects, as we consider these to be human-in-the-loop prompts, where the user must first read the paper to identify paper-specific aspects and then incorporate them into the prompt, which would require an experimental setup beyond our scope. This approach represents a form of human-AI collaboration, and we plan to explore this direction in future work.
@@
This work does not suggest or imply that the proposed task of predicting the aspects to focus on a given paper may replace human involvement. Instead, it is designed to assist reviewers by recommending aspects to consider. Reviewers retain full autonomy over their decisions regarding whether to include the suggested aspects. Since all the review data used in this study are publicly available and anonymized, processing it with commercial LLMs such as OpenAI GPT-4o does not raise ethical concerns.
@@ Acknowledgements
This work has been funded by the German Research Foundation (DFG) as part of the PEER project (grant GU 798/28-1), and funded/co-funded by the European Union (ERC, InterText, 101054961). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research).
@@ Appendix A Using the proposed aspect set to evaluate this paper
  • Related work: this paper reviews the opportunities and challenges of peer review in the era of LLMs, aspects in peer review, the quality of review writing, LLM-generated review detection, and scaffolding (Section 2).
@@
Appendix B More on aspect identification
@@ B.2 More on validity check
$<0.90$ Lack of definition, Context Definition, Paper

### v3

@@ Abstract
Peer review is central to academic publishing, but the growing volume of submissions is straining the process. This motivates the development of computational approaches to support peer review. While each review is tailored to a specific paper, reviewers often make assessments according to certain aspects such as Novelty, which reflect the values of the research community. This alignment creates opportunities for standardizing the reviewing process, improving quality control, and enabling computational support. While prior work has demonstrated the potential of aspect analysis for peer review assistance, the notion of aspect remains poorly formalized. Existing approaches often derive aspects from review forms and guidelines, yet data-driven methods for aspect identification are underexplored. To address this gap, our work takes a bottom-up approach: we propose an operational definition of aspect and develop a data-driven schema for deriving aspects from a corpus of peer reviews. We introduce a dataset of peer reviews augmented with aspects and show how it can be used for community-level review analysis. We further show how the choice of aspects can impact downstream applications, such as LLM-generated review detection. Our results lay a foundation for a principled and data-driven investigation of review aspects, and pave the path for new applications of NLP to support peer review. [^1] Our code and data are available at https://github.com/UKPLab/aspects-in-reviews.
@@ 1 Introduction
Peer review is an essential part of academic publishing. It is a complex, multifaceted process that requires a range of competencies including paper understanding, domain-specific knowledge, and critical thinking Shah (2022); Yuan et al. (2022). Ensuring review quality, especially among novice reviewers, is an open challenge Stelmakh et al. (2021a, b); Sun et al. (2024a). The increasing volume of publications puts further strain on the process, which motivates the development of computational approaches to support different stages of peer review, from reading the paper to the final decision-making by the program committees Arous et al. (2021); Checco et al. (2021); Stelmakh et al. (2021a); Shah (2022); Schulz et al. (2022); Yuan et al. (2022); Lin et al. (2023b); Kuznetsov et al. (2024).
Peer reviews are the central component of the reviewing process. While individual reviews can vary widely, reviewers within the same community tend to focus on a specific set of general quality categories, or aspects, such as Clarity and Novelty. These aspects can be found in review forms, instructional materials, guidelines, and the resulting review texts. Aspects allow comparison of submissions across different dimensions of quality. A comprehensive set of aspects shared among reviewers is critical for ensuring reviewing quality and consistency, and prior work has demonstrated the potential of aspect-based tools to support the review writing process Sun et al. (2024a, b).
Yet, several open questions remain. First, what is an aspect? While most prior work derives aspects from review forms and guidelines, the lack of an operational definition of aspect prevents the comparison of aspect schemata across studies. Second, what aspect granularity is appropriate for different tasks? While prior work operates with top-down, coarse-grained aspect schemata derived from review forms and guidelines, it is unclear whether they are comprehensive or provide sufficient granularity for NLP applications. Third, how can fine-grained aspect analysis support peer review? While aspects have been applied to certain tasks, the tasks that require a higher level of granularity are underexplored.
@@
Our work paves the path for data-driven analysis of aspects and enables new NLP applications to support peer review. Our method offers a bottom-up perspective that complements existing top-down schemata, and can be integrated with them by, for example, allowing domain experts to refine LLM-derived aspects into higher-level, domain-specific categories. Together, these perspectives enable a comprehensive aspect-based analysis of peer reviews. To summarize, we contribute the following:
@@ 2.2 Aspect in peer review
Beyond their role in review forms and guidelines, aspects are used in several other contexts related to peer review. Aspects have been applied to analyze the sentiment of reviews and discover the discourse relations within reviews Chakraborty et al. (2020); Kennard et al. (2022). In automatic review generation, Wang et al. (2020) generate reviews using aspect-based knowledge graphs, and Gao et al. (2024) prompt LLMs with aspect-based questions to generate reviews. Aspects are included in supporting systems for review writing, which have been shown to improve the comprehensiveness of the reviews written by reviewers with varying levels of experience Sun et al. (2024a, b). Table 1 summarizes the aspects used in these studies.
@@ 2.3 Quality of review writing
Defining the desiderata for review writing is essential to assess its effectiveness Jefferson et al. (2002). However, the desiderata are often not well defined or operationalizable in terms of automatic measurement Kuznetsov et al. (2024). While using simple proxies such as “helpfulness” may seem like a straightforward way to assess review quality, prior work suggests that such evaluation can be biased in several ways – for example, an evaluator may be biased towards longer reviews, or those that recommend acceptance of their papers Wang et al. (2021); Goldberg et al. (2023).
@@ 2.4 LLM-generated review detection
The strong capabilities of state-of-the-art LLMs in text generation have led to the need for detectors to identify LLM-generated contents and to prevent potential misuse Clark et al. (2021); Gao et al. (2023); Wu et al. (2023). In peer review, LLMs pose a risk that reviewers may exploit LLMs to produce reviews entirely, which raises serious ethical concerns Yu et al. (2024). Current strategies for detecting LLM-generated reviews align with methods for detecting LLM-generated texts: (a) LLM-as-a-judge, which prompts LLMs to identify LLM-generated contents Zheng et al. (2023), (b) detection models, which are fine-tuned on both human and LLM-generated texts Guo et al. (2023), (c) reference-based methods, which compare the similarity between a candidate text and one generated by an LLM Gehrmann et al. (2019); Ippolito et al. (2020); Liang et al. (2024a); Yu et al. (2024). In this work, we provide an alternative approach, and demonstrate a comprehensive, fine-grained aspect set helps the detection of LLM-generated reviews.
@@ 3.2 Method
To establish a comprehensive set of aspects, we selected reviews from NLP and machine learning (ML) conferences across different time periods. We randomly selected 50 papers from each of the NLPeer Dycke et al. (2023) and EMNLP23222Publicly available through the OpenReview API. datasets. We used the keywords “natural” and “language” to filter NLP-related papers from ICLR. We randomly selected 50 NLP-related papers from each of the ICLR conferences from 2020 to 2024. We selected 350 papers in total, corresponding to 1094 reviews. We segmented the reviews into sentences using Punkt Bird and Loper (2004) and performed the identification at the sentence level.
@@
We note that our post-processing method may lead to unrelated terms being grouped together. For example, using Improvement as a keyword groups Performance Improvement and Improvement Recommendation (i.e., suggestions to improve paper quality) together. We manually checked the results to verify and correct inappropriate groupings.
@@
See Appendix B for more details regarding the settings and human effort.
@@ 3.4 Validity check
Since LLMs have been shown to be sensitive to prompts Chen et al. (2023); Ajith et al. (2024), we experimented with different prompts and temperature to assess the consistency of the model annotations. We used three consistency metrics: exact match, BERTScore similarity Zhang et al. (2020), and Jaccard similarity. We show in Table 10 that the consistency between annotations generated under different settings is moderate to strong. 45.50% and 67.14% of the annotations obtained using different prompts and temperature have BERTScore similarities greater than 0.9. The increase in exact matches between the raw annotations and those mapped to the coarse label set suggests that differences introduced by varying prompts and temperature do not noticeably affect the categorization, as most of them still fall within the same label category. The Jaccard similarities of aspects mapped to the coarse label set also indicate strong consistency. Table 11 and 12 help interpret these scores by showing examples of how pairs of texts and sets correspond to different BERTScore and Jaccard similarities.
@@ 4 Aspect prediction
[…]
@@ 4.1 Method
RAP differs from aspect set construction described in Section 3 in that it is implemented using a supervised approach. We utilized our curated data to train models to identify aspects within reviews. We segmented the reviews into sentences using NLTK.
@@ 4.2 Results
In general, models perform better on RAP, and the coarse label set yields higher performance across both tasks.
@@ 5.2 Review comparison
We observe that LLM-generated reviews show a higher degree of similarity to each other in terms of aspects. This suggests that LLM-generated reviews may be more generic than human-written ones, with the model tending to comment on similar sets of aspects across different papers. This could also imply that the quality of LLM-generated reviews is still lacking in terms of specificity.
@@ 5.3 LLM-generated review detection
[…]
@@ 5.4 Recommendations
Based on our results, we make the following recommendations on selecting aspect granularity for a given application. The coarse label set is more suitable for high-level analysis tasks, such as analyzing review focus across different tracks and venues (Section 5.1). The fine label set is more appropriate for tasks that require nuanced analysis, where capturing specific and detailed feedback is critical, such as review comparison (Section 5.2) and LLM-generated review detection (Section 5.3). In some cases, a hybrid strategy might be the best option. For example, when evaluating review quality, one can first apply the coarse label set to assess coverage (i.e., whether key aspects are addressed), and then use the fine label set to assess specificity of the review. Aspect granularity is a design choice, and the optimal configuration depends on the task and user needs.
@@ 6 Conclusion
In this work, we focus on the NLP domain. While a comprehensive cross-domain analysis would be of great interest, it would require reviewing data and domain experts from other fields to construct a new taxonomy. As this would require a substantial study deserving a paper of its own, we leave it as potential future work.
@@
The LLM-generated annotations in this work are based on GPT-4o. While a comparative analysis across annotations generated by different models would be insightful, such a study would require not only applying different models, but also constructing multiple taxonomies, which needs expert involvement. Though valuable, this is beyond the scope of a single study.
For human annotation, we recognize that using aspect labels from our taxonomy may introduce potential label bias. A straightforward solution would be to ask the annotators to do an open-ended annotation without pre-defined labels. However, such study is very challenging in terms of the cognitive burden on annotators and ensuring annotation consistency. Automation bias is also a concern. While we address this by asking annotators to provide explanations for their annotations, follow-up work can explore alternative experimental strategies to measure and mitigate automation bias.
For the purposes of this study, we have kept the design of the PAP straightforward, treating it as a binary classification task predicting which aspects should be focused on given a paper. Yet, aspect relevance might indeed not be a binary decision, and modeling PAP as a ranking or regression problem could better reflect real world scenarios, where aspects have different levels of importance. Given the scope of the paper, we leave this investigation to future work.
@@
In Section 5.3, we only focus on the commonly used end-to-end general prompts for our review generator. We did not consider prompts that involve paper-specific aspects, as we consider these to be human-in-the-loop prompts, where the user must first read the paper to identify paper-specific aspects and then incorporate them into the prompt, which would require an experimental setup beyond our scope. This approach represents a form of human-AI collaboration, and we plan to explore this direction in future work.
@@
This work does not suggest or imply that the proposed task of predicting the aspects to focus on a given paper may replace human involvement. Instead, it is designed to assist reviewers by recommending the aspects to consider during their review. Reviewers retain full autonomy over their decisions regarding whether to include the suggested aspects. Since all the review data used in this study are publicly available and anonymized, processing it with commercial LLMs does not raise ethical concerns.
@@ Acknowledgements
This work has been funded by the German Research Foundation (DFG) as part of the PEER project (grant GU 798/28-1), and funded/co-funded by the European Union (ERC, InterText, 101054961). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work has been funded by the LOEWE Distinguished Chair “Ubiquitous Knowledge Processing”, LOEWE initiative, Hesse, Germany (Grant Number: LOEWE/4a//519/05/00.002(0002)/81). We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research).
@@ Appendix A Using the proposed aspect set to evaluate this paper
  • Related work: this paper reviews the opportunities and challenges of peer review in the era of LLMs, aspects in peer review, the quality of review writing, and LLM-generated review detection (Section 2).
@@
Appendix B More on aspect set construction
@@
Our method involves human effort on:
  • Post-processing LLM annotations: this is a one-time operation that involves both automated and manual steps. The automated part categorizes LLM annotations using high frequency terms and removes low frequency terms by setting a frequency threshold. The manual part focuses on verifying the remaining terms. In our work, this manual verification took one expert approximately 10 hours.
  • Taxonomy construction: we construct a taxonomy based on post-processed annotations. Mapping raw aspects into a taxonomy that reflects domain-relevant evaluation dimensions (such as the one shown in Table 2) involves both domain expertise and expert judgment. In our case, this step took one expert approximately 72 hours.
@@ B.2 More on validity check
<0.90<0.90 Lack of definition, Context Definition, Paper